OpenStudy (anonymous):

Among all pairs of numbers whose difference is 24, find a pair whose product is as small possible. What is the minimum product? I need all the steps please! in algebra form not calculus! i only have 1 hour>.<

OpenStudy (anonymous):

26-2?????-->26x2=52

OpenStudy (anonymous):

\[x - y = 24 and x = 24 + y , \therefore (24 + y)(y) = 24 y + y^2\]

OpenStudy (anonymous):

calc, but still.. should give u an indication

OpenStudy (anonymous):

\[x-y = 24\] \[xy = least\] \[xy = x if y = 1\] \[then x = \]

OpenStudy (anonymous):

x - y = 24 xy = least if xy = x,then xy = lest so numbers are 1,25